# Multi-Trigger Backdoor Attacks: More Triggers, More Threats

### v1

@@ Abstract
Backdoor attacks have emerged as a primary threat to (pre-)training and deployment of deep neural networks (DNNs). While backdoor attacks have been extensively studied in a body of works, most of them were focused on single-trigger attacks that poison a dataset using a single type of trigger. Arguably, real-world backdoor attacks can be much more complex, e.g., the existence of multiple adversaries for the same dataset if it is of high value. In this work, we investigate the practical threat of backdoor attacks under the setting of multi-trigger attacks where multiple adversaries leverage different types of triggers to poison the same dataset. By proposing and investigating three types of multi-trigger attacks, including parallel, sequential, and hybrid attacks, we provide a set of important understandings of the coexisting, overwriting, and cross-activating effects between different triggers on the same dataset. Moreover, we show that single-trigger attacks tend to cause overly optimistic views of the security of current defense techniques, as all examined defense methods struggle to defend against multi-trigger attacks. Finally, we create a multi-trigger backdoor poisoning dataset to help future evaluation of backdoor attacks and defenses. Although our work is purely empirical, we hope it can help steer backdoor research toward more realistic settings.
@@ 1 Introduction
Deep neural networks (DNNs) have become the mainstream models for computer vision (He et al., 2016; Dosovitskiy et al., 2021), natural language processing (Devlin et al., 2019; Mann et al., 2020), and speech recognition (Chan et al., 2016) tasks. However, studies have shown that DNNs are vulnerable to backdoor attacks (Gu et al., 2017; Chen et al., 2017; Li et al., 2020), which inject stealthy triggers into the victim (target) models during the training process by poisoning a small portion of the training data or manipulating the training procedure. In the context of image classification, the poisoning is typically done by adding a fixed but carefully designed trigger pattern to a few accessible (by the adversary) training images. The goal of a backdoor attack is to deceive the model into producing an adversary-specified output whenever the trigger pattern appears on a test input. With the wide adoption of large models pre-trained on unscrutinized web data or released by untrusted third parties, there has been an increasing concern about the backdoor vulnerability of these models, especially when deployed in safety-critical applications.
@@
A typical setup for backdoor attacks involves three key elements: 1) a predefined trigger pattern, 2) a few poisoning samples randomly selected from the victim dataset, and 3) a backdoor target label. Existing research efforts are mostly devoted to the design of effective and stealthy triggers like a fixed black-white pattern (Gu et al., 2017), adversarial noise (Turner et al., 2019), or sample-specific perturbations (Nguyen & Tran, 2020; Li et al., 2021a). In these attacks, there exists only one adversary and one single type of trigger (dataset-wise or sample-wise), thus we call them single-trigger attacks. This single-trigger setting is the only attack setting that has been studied extensively by existing works (Tran et al., 2018; Liu et al., 2018b, 2020; Barni et al., 2019; Nguyen & Tran, 2021). However, real-world attacks can be much more complex than single-trigger, raising the question of whether single-trigger attacks can pose practical threats to real-world applications. If attacking one particular dataset or AI system is of high value, then it should drive more than one adversary to launch an attack with different triggers. In this case, whether all triggers could coexist and be simultaneously effective becomes a concerning question.
In this paper, we extend the current backdoor research from single-trigger attacks to multi-trigger attacks [^1] We do not call them multi-adversary attacks because the triggers are our primary focus, and even one adversary could launch multiple attacks., where there exist multiple adversaries who simultaneously poison the same dataset using different types of triggers. Multi-trigger attacks can be implemented in three different ways: parallel, sequential, or in the form of hybrid-trigger that blends multiple trigger patterns into one super trigger pattern. Following the multi-trigger setup, we conduct a comprehensive analysis of 10 types of existing triggers and reveal the coexisting, overwriting, and cross-activating effects between these triggers. Specifically, we find that 1) different types of backdoor triggers can coexist (see Figure 1) as they can be injected into the same dataset either in parallel or sequentially, 2) certain triggers can overwrite other triggers at extremely low poisoning rates (e.g., $0.2\%$ or $0.5\%$), 3) one trigger can cross-activate other triggers under the sequential attack setting, and 4) different trigger patterns can be blended into one hybrid trigger to create more powerful attacks. These results reveal a greater security threat in the presence of multiple backdoor adversaries.
@@
Another concerning fact is that most existing backdoor defense methods were developed for single-trigger attacks, thus lacking proper evaluations against multi-trigger attacks. The missing of such evaluations could lead to overly optimistic views of backdoor robustness and a false sense of security. We find that unfortunately, it is indeed the case: multi-trigger backdoor attacks can greatly compromise current state-of-the-art backdoor detection and removal methods. Our main contributions can be summarized as follows:
  • We introduce the concept of Multi-Trigger Backdoor Attacks (MTBAs) to examine the practical threat of backdoor attacks in wild real-world scenarios where the same dataset or model could be attacked by multiple adversaries in parallel, sequentially, or via hybrid triggers.
  • By conducting more than 200 experiments with 10 types of triggers, 3 poisoning strategies (parallel, sequential, and hybrid-trigger), 4 poisoning rates ($[0.2\%-10\%]$), 3 label modification strategies (All2One, All2All, and All2Random), 2 datasets (CIFAR-10 and an ImageNet subset), 4 DNN architectures (2 CNNs and 2 ViTs), we reveal the co-existing, overwriting, and cross-activating effects of multi-trigger attacks.
  • By re-evaluating 8 existing backdoor defense methods (4 backdoor detection and 4 backdoor removal methods) against the parallel multi-trigger attack containing 10 types of triggers, we show that: 1) all detection methods struggle to detect All2All and All2Random attacks, albeit can detect All2One attacks to some extent; 2) none of the detection methods can detect the backdoor label of multi-trigger attacks; and 3) no backdoor removal methods can fully remove any of the 10 triggers. We also release a multi-trigger backdoored dataset to help evaluate future attacks and defenses.
@@ 2 Related Work
Existing backdoor defenses can be categorized into backdoor detection and backdoor removal (or mitigation) methods. Detection methods identify whether a given model has been backdoored by a backdoor attack (Chen et al., 2019; Wang et al., 2019; Guo et al., 2019; Xu et al., 2021) or whether a sample contains a backdoor trigger (Tran et al., 2018; Chen et al., 2019; Gao et al., 2019; Zeng et al., 2021; Huang et al., 2023). Removal methods aim to eliminate the backdoor trigger (if it exists) from the model without affecting its functionality. This can be done during the training process via anti-backdoor learning (Li et al., 2021b; Huang et al., 2022), or later on via fine-tuning, fine-pruning (Liu et al., 2018a), distillation (Li et al., 2021c), adversarial pruning (Wu & Wang, 2021), channel Lipschitzness based pruning (Zheng et al., 2022), or reconstructive pruning (Li et al., 2023). However, all of these defenses were developed for single-trigger attacks, thus facing high uncertainty when applied to defend multi-trigger attacks. In this paper, we run extensive experiments to reexamine the effectiveness of these defenses to multi-trigger backdoor attacks.
@@ 3.1 Threat Model and Problem Definition
In this work, we adopt a data poisoning and multi-trigger backdoor threat model. Specifically, there exist one or more adversaries where each adversary can access a few training samples but none of them can manipulate the training procedure. The adversaries can be independent or collusive, with independent adversaries poisoning different training samples with different types of triggers, while collusive adversaries could poison the same training samples in sequential order by stacking one trigger after another. There could also exist a super adversary that launches the above attacks all together by itself or combines multiple triggers into a hybrid and super trigger. We particularly focus on low to moderate poisoning rates ranging from 0.2% to 10%. The above settings of our threat model simulate the complex and wild real-world scenarios where multiple types of triggers can be injected into the same dataset by one or multiple adversaries under a small budget, for example, Carlini et al. managed to poison 0.01% of the LAION-400M (Schuhmann et al., 2021) or COYO-700M (Byeon et al., 2022) datasets for \$60 USD (Carlini et al., 2023).
@@
We call backdoor attacks that utilize multiple types of triggers (possibly from one or more adversaries) to attack the same dataset termed Multi-Trigger Backdoor Attacks (MTBAs). Figure 2 illustrates the idea of MTBA. Note that we did not call these attacks multi-adversary attacks as even one adversary could implement different types of triggers. The MTBA problem can also be formulated as Eq. (1), but with a slightly more complex poisoning set that contains multiple triggers:
@@ 3.2 Poisoning Strategies
MTBA is a straightforward extension of STBA. MTBA injects multiple types of triggers into the target model while STBA only injects one single type of trigger. Multiple backdoor triggers can be injected into the same dataset in different ways. Here, we formulate three poisoning strategies to simulate and represent the real-world scenarios: parallel, sequential, and hybrid-trigger poisonings.
@@ 3.2.2 Sequential Poisoning
Note that, although the adversaries follow a sequential order, we assume they are independent and their selected victim subsets do not overlap (${\mathcal{D}}_{s}^{k}\cap{\mathcal{D}}_{s}^{k+1}=\emptyset$) and only contain clean samples. These are reasonable assumptions because: 1) the adversaries have access to the victim samples, thus can easily ensure that they are clean (so would not impact trigger injection); and 2) independent attackers often own different samples. We also assume the subsequent adversary trains the backdoored model on the basis of the previous backdoored model. This is to simulate the current trend of fine-tuning large models, where a victim user may download and fine-tune a pre-trained (and poisoned by a previous adversary) large model for its own downstream application. For the same reason, we set the model training process to be as short as 10 epochs, i.e., Eq. (4) and Eq. (5) are optimized for only 10 epochs. Such a short optimization is enough to inject the trigger, while the overall training epochs are long enough for learning the clean task.
@@ 3.3 Label Modification
Following prior works (Gu et al., 2017; Xue et al., 2020), we explore the following three label modification strategies, including All2One, All2All, and All2Random (see Figure 3).
@@ 4.2.1 Parallel MTBAs
Refer to caption (a) All2One
Refer to caption (b) All2All
Refer to caption (c) All2Random
@@ 4.3 Hybrid-trigger MTBAs
Different from parallel and sequential MTBAs where the triggers are independent, our hybrid-trigger attack mixes multiple triggers into a single clean sample, all pointing to the same backdoor target label. This section explores the effectiveness of our hybrid-trigger attack on the CIFAR-10 dataset. To ensure high attack performance, we chose four triggers to construct the hybrid trigger, including one static trigger BadNets, one Dynamic, and two invisible triggers Nash and WaNet. Note that the Dynamic trigger does not work well as part of a hybrid. We apply those triggers in a random order (we will show that the order has minimum impact on the ASR) using a soft blending approach. The examples of hybrid-trigger are shown in appendix.

### v2

# Shortcuts Everywhere and Nowhere: Exploring Multi-Trigger Backdoor Attacks

@@ Abstract
Backdoor attacks have become a significant threat to the pre-training and deployment of deep neural networks (DNNs). Although numerous methods for detecting and mitigating backdoor attacks have been proposed, most rely on identifying and eliminating the “shortcut” created by the backdoor, which links a specific source class to a target class. However, these approaches can be easily circumvented by designing multiple backdoor triggers that create shortcuts everywhere and therefore nowhere specific. In this study, we explore the concept of Multi-Trigger Backdoor Attacks (MTBAs), where multiple adversaries leverage different types of triggers to poison the same dataset. By proposing and investigating three types of multi-trigger attacks including parallel, sequential, and hybrid attacks, we demonstrate that 1) multiple triggers can coexist, overwrite, or cross-activate one another, and 2) MTBAs easily break the prevalent shortcut assumption underlying most existing backdoor detection/removal methods, rendering them ineffective. Given the security risk posed by MTBAs, we have created a multi-trigger backdoor poisoning dataset to facilitate future research on detecting and mitigating these attacks, and we also discuss potential defense strategies against MTBAs.
@@ 1 Introduction
Deep neural networks (DNNs) have become the standard models for tasks in computer vision (He et al., 2016; Dosovitskiy et al., 2021), natural language processing (Devlin et al., 2019; Mann et al., 2020), and speech recognition (Chan et al., 2016). However, research has demonstrated that DNNs are susceptible to backdoor attacks (Gu et al., 2017; Chen et al., 2017; Li et al., 2020), where stealthy triggers are embedded into the target models during training by poisoning a small subset of the training data or altering the training process. In image classification, this typically involves adding a fixed, carefully crafted trigger pattern to a few training images that the adversary can access. The objective of a backdoor attack is to manipulate the model into producing a specific output chosen by the adversary whenever the trigger pattern appears in a test input. With the increasing use of large models pre-trained on unsupervised web data or those provided by untrusted sources, concerns about the backdoor vulnerability of these models have grown, particularly when they are used in safety-critical applications.
@@
Given the significant security threat posed by backdoor attacks on neural networks, a variety of backdoor detection and removal techniques have been developed. Some prominent methods include (Wang et al., 2019; Li et al., 2021c; Wu & Wang, 2021; Li et al., 2023). These methods generally operate under the assumption that backdoor triggers are unknown in practice. As a result, they focus on identifying backdoors by detecting the presence of “shortcuts” between a specific source class and a target class, and then mitigating the threat by eliminating these shortcuts. However, in this work, we systematically demonstrate that such approaches can be effectively bypassed by implementing multiple backdoor triggers that create shortcuts in all directions, rendering them indistinguishable and making it challenging to target any specific one for removal. This strategy exposes a critical vulnerability in existing backdoor defenses, highlighting the need for more robust and comprehensive protection mechanisms.
In this work, we extend traditional single-trigger backdoor attacks to multi-trigger backdoor attacks. Note that we do not call them multi-adversary attacks as our focus is on the triggers, demonstrating that even a single adversary can simultaneously employ multiple distinct triggers to poison a dataset. We explore three distinct implementation strategies for multi-trigger backdoor attacks: parallel, sequential, and hybrid, where the latter blends multiple trigger patterns into a single, more potent super trigger pattern. Furthermore, we conduct a thorough analysis of 10 different existing triggers, uncovering their coexistence, overwriting, and cross-activation effects. Our findings reveal that: 1) different backdoor triggers can coexist within the same dataset, whether injected in parallel or sequentially; 2) certain triggers can overwrite others even at minimal poisoning rates (as low as 0.2% or 0.5%); 3) one trigger can cross-activate another in sequential attack settings; and 4) different trigger patterns can be combined into a hybrid trigger to enhance attack efficacy. Figure 1 summarizes some of the experimental results. These insights highlight the heightened security risks posed by the presence of multiple backdoor adversaries.
@@
Next, we systematically evaluate existing backdoor defense methods since they have not been adequately evaluated against multi-trigger attacks. This gap in evaluation can lead to overly optimistic perceptions of backdoor robustness and a false sense of security. Our findings confirm this vulnerability: multi-trigger backdoor attacks can severely undermine the effectiveness of current state-of-the-art backdoor detection and removal techniques. Specifically, 1) widely used detection methods like Neural Cleanse (NC) (Wang et al., 2019) struggle to identify multi-trigger attacks, and 2) mainstream backdoor removal techniques such as fine-tuning (Liu et al., 2018a), Neural Attention Distillation (NAD) (Li et al., 2021c), Adversarial Neuron Pruning (ANP) (Wu & Wang, 2021), Reconstructive Neuron Pruning (RNP) (Li et al., 2023), and Anti-Backdoor Learning (ABL) fail to effectively purify models compromised by multiple triggers. The challenge of identifying multiple triggers (or the correct sequence to defend against them) is compounded by the difficulty of determining the exact number of triggers present, making defense against multi-trigger attacks extremely difficult.
Our main contributions are as follows.
  • We introduce the concept of multi-trigger backdoor attacks to expose the limitations of existing “shortcut”-based backdoor detection and removal methods. We highlight the practical threat posed by such attacks in real-world scenarios, where a dataset or model could be simultaneously attacked by multiple adversaries using parallel, sequential, or hybrid triggers.
  • Through more than 200 experiments involving 10 types of triggers, 3 poisoning strategies (parallel, sequential, and hybrid-trigger), 4 poisoning rates (ranging from 0.2% to 10%), 3 label modification strategies (All2One, All2All, and All2Random), 2 datasets (CIFAR-10 and an ImageNet subset), and 4 DNN architectures (including 2 CNNs and 2 ViTs), we uncover the coexistence, overwriting, and cross-activation effects of multi-trigger backdoor attacks.
  • By re-evaluating 8 existing backdoor defense methods (4 for detection and 4 for removal) against multi-trigger backdoor attacks with 10 types of triggers, we demonstrate that: 1) all detection methods struggle with All2All and All2Random attacks, though they show some effectiveness against All2One attacks; 2) none of the detection methods can accurately identify the target of multi-trigger attacks; and 3) no backdoor removal method is capable of fully eliminating any of the 10 triggers.
  • We have built and released a multi-trigger backdoor dataset to support future research on backdoor attacks and defenses.
@@ 2 Related Work
Existing backdoor defenses can be categorized into backdoor detection and backdoor removal (or mitigation) methods. Detection methods identify whether a given model has been backdoored by a backdoor attack (Chen et al., 2019; Wang et al., 2019; Guo et al., 2019; Xu et al., 2021) or whether a sample contains a backdoor trigger (Tran et al., 2018; Chen et al., 2019; Gao et al., 2019; Zeng et al., 2021; Huang et al., 2023).
Removal methods aim to eliminate the backdoor trigger (if it exists) from the model without affecting its functionality. This can be done during the training process via anti-backdoor learning (Li et al., 2021b; Huang et al., 2022), or later on via fine-tuning, fine-pruning (Liu et al., 2018a), distillation (Li et al., 2021c), adversarial pruning (Wu & Wang, 2021), channel Lipschitzness based pruning (Zheng et al., 2022), or reconstructive pruning (Li et al., 2023). However, all of these defenses were developed for single-trigger attacks, thus facing high uncertainty when applied to defend multi-trigger attacks. In this paper, we run extensive experiments to reexamine the effectiveness of these defenses to multi-trigger backdoor attacks.
@@ 3.1 Threat Model and Problem Definition
Our threat model introduces 3 poisoning strategies for multi-trigger attacks: parallel, sequential, and hybrid, covering scenarios involving both independent and collusive adversaries. Specifically, there exist one or more adversaries where each adversary can access a few training samples but none of them can manipulate the training procedure. The adversaries can be independent or collusive, with independent adversaries poisoning different training samples with different types of triggers, while collusive adversaries could poison the same training samples in sequential order by stacking one trigger after another. There could also exist a super adversary that launches the above attacks altogether or combines multiple triggers into a hybrid trigger.
@@
We call backdoor attacks that utilize multiple types of triggers (possibly from one or more adversaries) to attack the same dataset termed Multi-Trigger Backdoor Attacks (MTBAs). Figure LABEL:fig:overview illustrates the idea of MTBA. Note that we did not call these attacks multi-adversary attacks as even one adversary could implement different types of triggers. The MTBA problem can also be formulated as Eq. (1), but with a slightly more complex poisoning set that contains multiple triggers:
@@ 3.2 Poisoning Strategies
Conceptually, MTBA is an extension of STBA. While STBA injects only one type of trigger into the target model, MTBA introduces multiple types of triggers. We introduce three poisoning strategies to simulate diverse real-world threat scenarios: parallel, sequential, and hybrid-trigger.
@@ 3.2.2 Sequential Poisoning
Note that, although the adversaries follow a sequential order, we assume they are independent and their selected victim subsets do not overlap (${\mathcal{D}}_{s}^{k}\cap{\mathcal{D}}_{s}^{k+1}=\emptyset$) and only contain clean samples. These are reasonable assumptions because: 1) the adversaries have access to the victim samples, thus can easily ensure that they are clean (so would not impact trigger injection); and 2) independent attackers often own different samples. We also assume the subsequent adversary trains the backdoored model on the basis of the previous backdoored model. This is to simulate the current trend of fine-tuning large models, where a victim user may download and fine-tune a pre-trained (and poisoned by a previous adversary) large model for its own downstream application.
@@ 3.3 Label Modification
In the context of multi-trigger backdoor attacks, different adversaries may share the same backdoor target if it is of general interest, or they may have entirely different (or random) target labels. Following prior works (Gu et al., 2017; Xue et al., 2020), we explore the following three label modification strategies, including All2One, All2All, and All2Random (see Figure 3).
@@ 4.2.1 Parallel MTBAs
(a) All2One
(b) All2All
(c) All2Random
@@ 4.3 Hybrid-trigger MTBAs
Different from parallel and sequential MTBAs where the triggers are independent, our hybrid-trigger attack mixes multiple triggers into a single clean sample, all pointing to the same backdoor target label. Figure 6 shows example of the hybrid triggers. This section explores the effectiveness of our hybrid-trigger attack on the CIFAR-10 dataset. To ensure high attack performance, we chose four triggers to construct the hybrid trigger, including one static trigger BadNets, one Dynamic, and two invisible triggers Nash and WaNet. Note that the Dynamic trigger does not work well as part of a hybrid. We apply those triggers in a random order (we will show that the order has minimum impact on the ASR) using a soft blending approach. The examples of hybrid-trigger are shown in appendix.